OpenStudy (thatonegirl_):

Probability that x is at least 3? Probability of 1=.1 2=.3 3=.4 4=.2

OpenStudy (kirbykirby):

\(P(X \ge 3) = P(X=3)+P(X=4)\) So add the probabilities associated with X=3 and X=4